Suppose you held a diversified portfolio consisting of a $7,500 investment in each of 20 different common stocks. The portfolio's beta is 2.30. Now suppose you decided to sell one of the stocks in your portfolio with a beta of 1.0 for $7,500 and use the proceeds to buy another stock with a beta of 1.88. What would your portfolio's new beta be? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

Increase in beta as a result of net sales=(1.88-1)=0.88

Hence increase/stock=0.88/20=0.044

Hence new beta=2.3+0.044

=2.34(Approx).